In a house, the rectangular living room has a perimeter of 96ft. The length is 8ft less than twice the width. Express this information in the form of two equations.
Answer by checkley77(12844)

L=2W-8
2L+2W=96
REPLACE L BY (2W-8) IN THE SECOND EQUATION & SOLVE FOR W.
2(2W-8)+2W=96
4W-16+2W=96
6W=96+16
6W=112
W=112/6
W=18.667 ANS. FOR THE WIDTH.
L=2*18.667-8
L=37.333-8
L=29.333 ANS. FOR THE LENGTH.
PROOF:
2*29.333+2*18.667=96
58.667+37.333=96
96=96